OpenStudy (anonymous):

lim 1/(x+2)^2 x-> -2 from the left

OpenStudy (anonymous):

\[\lim_{x \rightarrow -2-} 1/(x+2)^2 \]

OpenStudy (anonymous):

denominator goes to zero through positive numbers (because it is a square) so the limit is \(\infty\)

OpenStudy (anonymous):

can you show work like steps because im confused

OpenStudy (anonymous):

i cannot think of any steps to show the numerator is constantly 1, and the denominator is getting closer and closer to zero

OpenStudy (anonymous):

the only possible question is, does it go to \(-\infty\) or \(\infty\) and it is the latter since \((x+2)^2>0\)

OpenStudy (anonymous):

how do you know its greater then 0?

OpenStudy (anonymous):

because any number squared is positive